### v10

# On the representation-free formalism in quantum theory

Abstract
The widespread bra-ket formalism offers valuable tools for conducting representation-free considerations in quantum theory. However, it is not without its drawbacks. In this work, we discuss these drawbacks in detail and subsequently construct a new representation-free scheme. Similar to the bra-ket formalism, the present scheme provides essential means to facilitate representation-free considerations. At the same time, it is entirely free from the drawbacks of the bra-ket formalism. Unlike the dual-space bra-ket formalism, the present scheme allows both one-space and dual-space interpretations, which is a beneficial feature. The constructed universal scheme is very well-suited for performing representation-free practical calculations.
pacs:
03.65.-w, 03.65.Ca, 03.65.Ta
I Introduction
In quantum theory, the representation-free approach is indispensable when working with general relations. It seems appropriate to compare the strengths of this approach with those of, for example, the vector form of the Maxwell equations. Typically, physicists adopt the bra-ket formalism as the representation-free approach. This formalism provides helpful tools dir1 ; dir2 for conducting practical calculations. However, it is not satisfactory to a substantial degree, and various authors drew attention to its drawbacks gie ; ja ; gr ; wei ; si .
Besides those drawbacks, the bra-ket formalism is a dual-space scheme, which presents a substantial limitation. A more straightforward and natural one-space formulation would be preferable in many cases. Mathematicians typically employ the formulation, listed below, of such a type in studies both on abstract Hilbert space and quantum theory. Weinberg wei also utilized this formulation. This formulation does not have drawbacks of the bra-ket formalism but it lacks helpful tools that the bra-ket formalism provides. Furthermore, while this one-space formulation does not include bra vectors, they are still required for various applications, and the dual-space approach incorporating bra vectors is needed for noteworthy purposes, see, e.g., Refs. gie ; gel ; bo ; oh ; oh1 ; gan ; ott . Until now, the bra-ket formulation and this one-space formulation have been the only ones used in the literature.
In Sec. 3, we construct a convenient and universal representation-free scheme that, in contrast to the aforementioned formulations, allows both the one-space and dual-space interpretations. Furthermore, unlike the aforementioned mathematically oriented formulation, the present scheme provides tools that facilitate practical calculations, of the type of tools inherent in the bra-ket formalism. With all that, the present formulation is entirely free from the drawbacks of the bra-ket formalism.
Before this, we address the strengths and drawbacks of the bra-ket formalism in Sec. 2. In particular, the original version dir1 of this formalism is critically analyzed in detail, which, to our knowledge, has not been done previously.
One may note that any representation-free formulation of quantum theory should be equivalent to others, provided that they all are correct and complete. However, this by no means implies that they would be equally suitable for practical calculations.
II Bra-ket formalism
In the first part of this section, we will analyze the original version of the bra-ket formalism given in the classical book dir1 . There exists a later version of this formalism discussed in Ref. gie , which we will describe next. Following that, we will outline the appealing properties of the bra-ket formalism, and discuss the drawbacks of its later version.
The present analysis of the original version of the bra-ket formalism is needful, considering that this version is still followed either in full, as in Ref. mes , or ”in part,” as in Ref. co . A clear distinction between the two versions has not been made in the literature. [^2] In Ref co , the reasoning from Ref. dir1 is reproduced regarding the action of operators on bra vectors and the definition of adjoint operators. However, the definition of bra vectors themselves adopted there, which is the same as in Subsection B, differs from that in Ref. dir1 and is incompatible with the aforementioned reasoning. In the considerations of this section, state vectors are assumed to have finite norms.
II.1 Original version
According to Secs. 6-8 of Ref. dir1 , the bra-ket scheme is constructed through the following main steps.
1. One considers linear functionals $F$ that depend on state vectors $|v\rangle$. One assumes that there exists an anti-linear one-to-one correspondence between the set of these functionals and the set of state vectors themselves. Denoting state vectors as $|u\rangle$ in this context, this correspondence can be expressed as $F\Leftrightarrow|u\rangle$. To designate the functionals $F$ based on this correspondence, the symbol $\langle u|$ is introduced: $F\Rightarrow\langle u|$. Correspondingly, the values of the functionals $F(|v\rangle)$ are denoted as $\langle u|v\rangle\equiv\langle u|(|v\rangle)$. The quantities $\langle u|v\rangle$, referred to as scalar products, are thus linear in $|v\rangle$ and anti-linear in $|u\rangle$. It is assumed that $\langle u|v\rangle^{*}=\langle v|u\rangle$ which is consistent with the above. It is also assumed that $\langle v|v\rangle\geq 0$ for any $|v\rangle$.
The functionals $\langle u|$ are referred to as bra vectors. State vectors are called ket vectors. The ket and bra vectors that form a $\langle u|\Leftrightarrow|u\rangle$ pair are termed conjugates of each other.
2. The action of linear operators on bra vectors is defined. Let $O$ be a linear operator that acts on ket vectors. To give meaning to its action on bra vectors, the scalar products of the form $\langle u|\bigl{(}O|v\rangle\bigr{)}$ are considered. These scalar products are viewed, for a given bra vector $\langle u|$ and a given operator $O$, as the values of a linear functional that depends on states $|v\rangle$. According to the definition above, this functional is some bra vector. Denoting this bra vector as $\langle w|$, one then may express the values of this functional as $\langle w|v\rangle$, as above. That is $\langle u|\bigl{(}O|v\rangle\bigr{)}=\langle w|v\rangle$.
Further, for a given operator $O$, the family of the $\langle u|\bigl{(}O|v\rangle\bigr{)}$ functionals specified by the bra vectors $\langle u|$ is considered. The operator $O$, acting in the space of ket vectors, uniquely defines an operator in the space of bra vectors that realizes the mapping $\langle u|\Rightarrow\langle w|$ with the above definition of $\langle w|$. Then, one designates the latter operator with the same symbol $O$: $\langle w|=\langle u|O$. With this definition, one has
$\bigl{(}\langle u|O\bigr{)}|v\rangle=\langle u|\bigl{(}O|v\rangle\bigr{)}\stackrel{{\scriptstyle def}}{{=}}\langle u|O|v\rangle.$ (1)
The operator in the space of bra vectors, thus introduced, is linear.
3. The adjoint operator $O^{\dagger}$ of a linear operator $O$ is defined. The definition consists of the statement that the ket vector $O^{\dagger}|u\rangle$ is conjugate of the bra vector $\langle u|O$ defined above. That is,
$\langle v|\bigl{(}O^{\dagger}|u\rangle\bigr{)}\stackrel{{\scriptstyle def}}{{=}}\left[\bigl{(}\langle u|O\bigr{)}|v\rangle\right]^{*}.$ (2)
Using this definition along with Eq. (1) applied to both $O^{\dagger}$ and $O$ operators, one then gets the relation
$\langle v|\bigl{(}O^{\dagger}|u\rangle\bigr{)}=\bigl{(}\langle v|O^{\dagger}\bigr{)}|u\rangle\stackrel{{\scriptstyle def}}{{=}}\langle v|O^{\dagger}|u\rangle=\langle u|O|v\rangle^{*}.$ (3)
Let us discuss the listed reasoning. First, consider step 1 above. We shall need the following definition. A linear functional $F$ is said to be unbounded if states $|v\rangle$ can be found such that the quantity $\langle v|v\rangle^{-1/2}|F(|v\rangle)|$ exceeds any given constant. Otherwise, $F$ is said to be bounded. Jauch ja has noted that it is incorrect to define bra vectors in terms of unbounded linear functionals. Here, we shall give a simple explanation of this. In our case, $F=\langle u|$, and $F(|v\rangle)=\langle u|v\rangle$. According to the Schwarz inequality (that follows merely from the definition of the scalar product, see, e.g., reed ), one has
$\langle v|v\rangle^{-1/2}|\langle u|v\rangle|\leq\langle u|u\rangle^{1/2}.$
The right-hand side of this inequality is a constant independent of $|v\rangle$. Thus, the linear functional $\langle u|$ cannot be unbounded. That is, not any linear functional $F=\langle u|$ can be put into correspondence with a state vector $|u\rangle$ in a way that $F(|v\rangle)=\langle u|v\rangle$ possesses the prescribed properties of the scalar product. In other words, the definition of bra vectors in the reasoning listed above is incompatible with the natural properties of the scalar product that are adopted within that reasoning. Thus, in contradiction to that reasoning, not every linear functionals $F(|v\rangle)$ possesses the prescribed properties of bra vectors. Unbounded $F(|v\rangle)$ do not exhibit these properties. (Certain bounded $F(|v\rangle)$ are also not bra vectors. These include some of $F(|v\rangle)$ involved in the step 2 of the aforementioned reasoning, as can be concluded from the considerations that follow.)
Jauch ja has suggested associating bra vectors exclusively with bounded linear functionals (see also below). We contend, however, that this is insufficient to cure the formalism. The point is as follows. The reasoning in step 2 above, which addresses the action of operators on bra vectors, was based on the assertion that the linear functionals $F$ of the form $F(|v\rangle)=\langle u|\bigl{(}O|v\rangle\bigr{)}$ can be considered bra vectors. If one denotes these bra vectors as $\langle w|$, the relation expressing this assertion can be written as $\langle u|\bigl{(}O|v\rangle\bigr{)}=\langle w|v\rangle$. ($\langle w|\Rightarrow\langle u|O$.) Let us show that this relation is, in general, invalid since the state $|w\rangle$ may not exist.
For example, let us consider a self-adjoint operator $O$ with a discrete unbounded spectrum, specifically focusing on the single-particle problem and the operator ${\hat{l}}^{2}$ of the square of orbital momentum. Let $|nlm\rangle$ be eigenstates of the ${\hat{l}}^{2}$ operator, ${\hat{l}}^{2}|nlm\rangle=l(l+1)|nlm\rangle$. Normalize them to unity. In the relation we are discussing, with $O={\hat{l}}^{2}$, let us take $|nlm\rangle$ as $|v\rangle$. Then this relation transforms to $l(l+1)u^{*}(nlm)=w^{*}(nlm)$ where $u(nlm)=\langle nlm|u\rangle$ and $w(nlm)=\langle nlm|w\rangle$ are the wave functions of the states $|u\rangle$ and $|w\rangle$ in the nlmnlm representation.
Suppose, for example, that $|u\rangle$ is a state with fixed $n$ and $m$ values. Assuming that the state $|w\rangle$ entering the relation under discussion exists, we can calculate the norm of the state $|w\rangle-\sum_{l\geq|m|}^{l_{max}}w(nlm)|nlm\rangle$. This leads to the usual-type inequality: $\sum_{l\geq|m|}^{l_{max}}|w(nlm)|^{2}\leq\langle w|w\rangle$. From this inequality, it follows that the series $\sum_{l\geq|m|}^{\infty}[l(l+1)]^{2}|u(nlm)|^{2}$ must be convergent for the state $|w\rangle$ to exist. However, it is clear that this condition is not universally satisfied, as the only requirement regarding the state $|u\rangle$ is that the series $\sum_{l\geq|m|}^{\infty}|u(nlm)|^{2}$ is convergent. Thus, for certain states $|u\rangle$ and operators $O$ of interest, the equality of the form $\langle u|\bigl{(}O|v\rangle\bigr{)}=\langle w|v\rangle$ cannot hold. Consequently, the reasoning of step 2 we are discussing is not valid. This implies that, in addition to bra vectors themselves, the action of operators on bra vectors also cannot be defined according to this reasoning.
Let us also note that, even with the sum $\sum_{l\geq|m|}^{\infty}|u(nlm)|^{2}$ being convergent, the behavior $l(l+1)|u(nlm)|\rightarrow\infty$ at $l\rightarrow\infty$ is still possible. At least for states $|u\rangle$ exhibiting this behavior, the functionals $F(|v\rangle)=\langle u|\bigl{(}{\hat{l}}^{2}|v\rangle\bigr{)}$ are unbounded and cannot be presented in the form $\langle w|v\rangle$ for this reason alone.
And, as for the definition of the adjoint operator, which corresponds to step 3 of the reasoning under discussion, this definition involves objects like $\langle u|O$. However, as argued above, such objects are, in fact, not properly defined in the previous step of the reasoning. Therefore, the definition of the adjoint operator being discussed also cannot be adopted. A different definition is listed below.
II.2 Revised version
[…]
Let us now list the revised definitions of the action of an operator on a bra vector and of the adjoint operator, see, e.g., Ref. gie . In contrast to the previously listed version of the bra-ket formalism, now the adjoint operator is introduced first. It is defined in accordance with the relation
$\langle v|\bigl{(}O^{\dagger}|u\rangle\bigr{)}=\left[\langle u|\bigl{(}O|v\rangle\bigr{)}\right]^{*}$ (4)
Eq. (4) is different from Eq. (2).
The revised definition for the action of an $O$ operator on a covector is as follows: the covector $\langle u|O$ is defined as the conjugate of the vector $O^{\dagger}|u\rangle$ where the operator $O^{\dagger}$ is given by Eq. (4). With this definition, the equality (1) is valid due to the relation (4). The definition for the action of an $O^{\dagger}$ operator on a covector is that the covector $\langle v|O^{\dagger}$ is the conjugate of the vector $O|v\rangle$. With this definition, the equalities (3) are valid due to the relation (4).
II.3 Features of the bra-ket scheme
Efficient ”working tools” provided by the bra-ket formalism dir1 include a special designation $|\,\,\rangle$ for marking state vectors, an ”explicit” form of the projection-type operators, such as $|u\rangle\langle v|$, and a convenient explicit representation of eigenstates of observables that form a complete set, such as $|{\bf p}_{1}s_{z1},\ldots,{\bf p}_{n}s_{zn}\rangle$ (where ${\bf p}_{i}$ and $s_{zi}$ denote the momenta and projections of spins of particles). Matrix elements of operators between states belonging to a basis set then take a convenient form, such as $\langle n_{1}^{\prime},n_{2}^{\prime},\ldots,n_{k}^{\prime}|O|n_{1},n_{2},\ldots,n_{k}\rangle$. (In our scheme described below, all such tools are also available.)
Further, one needs to consider the drawbacks of the above-described revised version of the bra-ket scheme. First, let us examine the conventional bra-ket expressions for matrix elements, such as $\langle u|O|v\rangle$. The reasoning on this subject in this paragraph is basically similar to the considerations by Gieres gie . The expressions $\langle u|O|v\rangle$ is introduced through the equalities (1). However, these equalities are not universally valid. Indeed, let $D(O)$ and $D(O^{\dagger})$ denote the domains of definition of the operators $O$ and $O^{\dagger}$. The hermiticity relation (4), on which the equalities (1) are based, applies only when $|v\rangle\in D(O)$ and $|u\rangle\in D(O^{\dagger})$ simultaneously. To provide a complete definition of the $O^{\dagger}$ operator, the domain $D(O^{\dagger})$ is to be specified. It is to be accepted that the (maximal) domain $D(O^{\dagger})$ consists of all the $u$ vectors for which the relation (4) takes place for any $v$ belonging to $D(O)$. Suppose that $u\notin D(O^{\dagger})$. In this case, the above definition of the covector $\langle u|O$ is unapplicable, and the matrix element $\bigl{(}\langle u|O\bigr{)}|v\rangle$ entering Eq. (1) does not exist. Suppose now that $v\notin D(O)$. Then the matrix element $\langle u|\bigl{(}O|v\rangle\bigr{)}$ that appears in Eq. (1) does not exist. In both cases, the expression $\langle u|O|v\rangle$ is undefined. Nevertheless, the matrix elements $(u,Ov)$ can exist even when $u\notin D(O^{\dagger})$ (assuming $D(O^{\dagger})\neq{\cal H}$) and the matrix elements $(O^{\dagger}u,v)$ can exist even when $v\notin D(O)$ (assuming $D(O)\neq{\cal H}$). In such cases, the well-defined matrix elements $(u,Ov)$ and $(O^{\dagger}u,v)$ cannot be represented by the conventional bra-ket expression $\langle u|O|v\rangle$.
One cannot disregard such matrix elements $(u,Ov)$ and $(O^{\dagger}u,v)$. Consider the case of operators $O$ that can correspond to observable quantities, specifically self-adjoint operators with unbounded spectra. In this case, the expression $\langle u|O|v\rangle$ is well-defined only if $|u\rangle\in D(O)$ and $|v\rangle\in D(O)$ simultaneously. However, for any operator $O$ of this class, there exist, say, states $|v\rangle$ that do not belong to $D(O)$. Indeed, when the spectrum of $O$ is discrete, writing $\bigl{(}\langle v|O\bigr{)}\bigl{(}O|v\rangle\bigr{)}=\sum_{\lambda,i}|\lambda\langle\lambda,i|v\rangle|^{2}$, where $\lambda$ and $|\lambda,i\rangle$ are the eigenvalues and eigenstates of the operator $O$, one sees that there always exist states $|v\rangle$ for which the series diverges, indicating that the state $O|v\rangle$ does not exist. When the spectrum of $O$ is continuous, a similar argument applies. [^4] The more general Hellinger-Töplitz theorem (see, e.g., reed ) also states that $D(O)\neq{\cal H}$ always holds.
Thus, matrix elements $(Ou,v)$ that cannot be represented in the form $\langle u|O|v\rangle$ always exist for the aforementioned operators $O$. For example, let $p_{z}$ be the operator of a component of momentum. Suppose that for some $n$, the state $p_{z}^{n}|u\rangle$ is well-defined while the state $p_{z}^{n}|v\rangle$ is not well-defined. That is, the $u({\bf p})$ wave function decreases sufficiently fast as $p_{z}$ increases, while the $v({\bf p})$ wave function does not. In this case, the well-defined matrix element $\bigl{(}p_{z}^{n}u,v\bigr{)}$ cannot be represented in the bra-ket form $\langle u|p_{z}^{n}|v\rangle$.
Thus, instead of conventional bra-ket matrix elements of the $\langle u|O|v\rangle$ type, one, in general, should have used expressions like $\langle u|\bigl{(}O|v\rangle\bigr{)}$, $\bigl{(}\langle u|O\bigr{)}|v\rangle$, or $\bigl{(}\langle u|O\bigr{)}\bigl{(}O^{\prime}|v\rangle\bigr{)}$, which is overly cumbersome.
It was suggested gie to adopt the rule that in expressions like $\langle u|O|v\rangle$, the operator acts only to the right. However, this rule does not extend to matrix elements such as $\bigl{(}\langle u|O\bigr{)}\bigl{(}O^{\prime}|v\rangle\bigr{)}$. Furthermore, for expressions of the $\bigl{(}\langle u|O\bigr{)}|v\rangle$ type, this rule is inconvenient, as it requires replacing them with their complex conjugate each time one encounters them.
Another drawback of the bra-ket scheme is that covectors of the form $\langle u|O$ entail potential ambiguities regarding the domain of definition gie . To avoid this inconvenience, it has been suggested gie ; co to represent the covector $\langle u|O$ with the expression $\langle O^{\dagger}u|$ implying that $u\equiv|u\rangle$ in this context. However, in this case, the use of explicit expressions for vectors specified by quantum numbers is precluded.
Herewith, we note that while the bra-ket formalism incorporates adjoint operators as a constituent, it is possible to proceed without adjoint operators entirely in this context, as demonstrated in the next section. [^5] It is worth mentioning that, although this rarely occurs in practice, there may exist more than one operator that is adjoint to a given operator, as F. Gieres mentioned to me. Therefore, relying on the one-to-one correspondence between an operator and its adjoint inherent in the bra-ket formalism is, in principle, also a limitation.
In addition to all the issues mentioned above, a significant disadvantage of the bra-ket formalism is that this formalism, being a dual-space formalim, thus excludes the natural one-space point of view.
There are known serious difficulties that students encounter with the bra-ket formalism. In Ref. si , for instance, the authors write in the Abstract: ”We have been investigating the difficulties that the advanced undergraduate and graduate students have with Dirac notation.…We find that many students struggle with Dirac notation, and they are not consistent in using this notation”. Probably, these difficulties might be alleviated if the one-space picture were employed to realize the representation-free approach. This picture would highlight the similarity of this approach to simple vector algebra in its conventional one-space form, as it is permanently utilized throughout all branches of physics.
Even though the bra-ket scheme is typically adopted as a representation-free formalism, it has faced criticisms in the literature. Gieres gie characterizes the bra-ket scheme as ”representing a purely symbolic calculus”. An utterance gr , also cited in gie , states: “Unfortunately, the elegance, outward clarity and strength of Dirac’s formalism are gained at the expense of introducing mathematical fictions.…One has a formal ‘machinery’ whose significance is impenetrable, especially for the beginner, and whose problematics cannot be recognized by him”. Also Steven Weinberg wei expresses his refusal to use the bra-ket formalism speaking of its drawbacks. Instead, he employs the aforementioned representation-free, mathematically oriented formulation, while acknowledging that it also has its disadvantages.
III Present formulation
First, we would like to justify the necessity for a new formulation of the representation free approach concisely. The mathematically oriented one-space formulation of this approach lacks practical calculation tools, such as those provided by the dual-space bra-ket formalism. Conversely, the latter formalism also has its drawbacks considered above. Gieres gie suggested that the representation-free approach itself is not that useful because the bra-ket formalism does not adequately address the domains of definitions of operators. Furthermore, he argued that one should apply the coordinate representation to a greater extent, as these domains of definitions are naturally determined in this representation.
However, due to the significant advantages of the representation-free approach when working with general relations, a much better way out would be to construct a realization of this approach that is more suitable than the bra-ket formalism in the aforementioned and other respects. While, to obtain information regarding the specific domains of definitions of the involved operators, one could utilize suitable representations before transitioning to a representation-free description.
An appropriate realization of the representation-free approach should incorporate all the tools of the type of those in Dirac’s formalism, that facilitate practical calculations. It should also utilize a one-space description while allowing for a dual-space interpretation simultaneously. The dual-space description is needed for specific noteworthy purposes, such as rigged Hilbert space gel applications gie ; bo ; oh ; oh1 or refined interpretations of quantum theory gan ; ott . Additionally, a dual-space interpretation would align more closely with the widely used bra-ket formalism. The representation-free scheme constructed here meets all these requirements.
Let us move on to the present representation-free scheme. As the bra-ket formalism, the present scheme includes a special marking of state vectors. We denote state vectors specified by quantum numbers as ${\bf p}_{1}s_{z1},\ldots,{\bf p}_{n}s_{zn}/$ (and, for stationary scattering states tay , as ${\bf p}_{1}s_{z1},{\bf p}_{2}s_{z2}\pm/$). We also mark state vectors as /u//u/ in certain other cases, but we simply use symbols like $u$ when this does not lead to confusion.
For the quantum mechanical scalar product, we suggest the expression like $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ or, for instance, ${\bf r}/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}/plm/$, or ${\bf p}_{1},\ldots,{\bf p}_{n}/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$. While in the bra-ket scalar product expression $\langle u|v\rangle$ the bra vector $\langle u|$ is a linear functional, in the present $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ expression $u$ is a vector. Thus, $u$ and $v$ belong to the same space. The $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ expression is an analog to the expression for the scalar product of vectors in finite-dimensional spaces, such as ${\bf u}\cdot{\bf v}$. (Of course, $\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}u=(u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v)^{*}, vv$$\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v\geq 0$, and $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ is linear in $v$ and thus anti-inear in $u$.) The adjoint operator $O^{\dagger}$ of a linear operator $O$ is defined in the conventional way:
$O^{\dagger}u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v=u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}Ov$ (5)
where the domain of definition of $O^{\dagger}$ is indicated in relation to Eq. (4). (If $O$ is anti-linear then $O^{\dagger}u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v=Ov\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}u$.) Eq. (5) is simpler than Eq. (4). Unlike the bra-ket formalism, adjoint operators are not a component of the present scheme. For this reason and others, this scheme is less involved than the bra-ket one.
[…]
In some (though not all) cases, it is convenient to omit the dot in the scalar products of the form $u/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}O/v/$ writing them simply as /u/O/v//u/O/v/. (One has $u/O/v/^{*}=O/v/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}/u/$.) Matrix elements of an operator in a basis are then represented like
$n_{1}^{\prime},n_{2}^{\prime},\ldots n_{k}^{\prime}/O/n_{1},n_{2},\ldots n_{k}/.$
In the case of matrix elements $\ldots j_{1}m_{1}/O_{jm}/\ldots j_{2}m_{2}/$ of an irreducible tensor operator $O_{jm}$ between states with specified angular momenta, it is natural to denote the corresponding reduced matrix elements as $\ldots j_{1}//O_{j}//\ldots j_{2}/$.
We introduce multiplication signs * between constituents of emerging expressions, such as scalar products or state vectors and scalar products. This appears as $u*v\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}w$, or $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v*u^{\prime}\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v^{\prime}*w$. The multiplication signs provide clarity, and in particular, they are required for what follows.
One should distinguish between the expressions $cu\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ and $(cu)\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$, where $c$ is a constant. While the second expression represents the scalar product of the states $cu$ and $v$, the first one is interpreted as the product of the constant $c$ and the scalar product $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ (unlike expressions such as $Ou\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$). One has $(\sum_{i}c_{i}u_{i})\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v=\sum_{i}(c_{i}u_{i})\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v=\sum_{i}c_{i}^{*}u_{i}\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$. If, for example, in the above relations the constants $c$ or $c_{i}$ themselves are scalar products then for clarity the replacement $c$ by cc* or $c_{i}$ by $c_{i}*$ and $c_{i}^{*}$ by $c_{i}^{*}*$ is to be done.
Further, as in the case of the bra-ket formalism, we propose an ’explicit’ expression for projection-type operators. In the present scheme, this expression is of the form $u*v\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. It incorporates the multiplication sign between the vectors $u$ and $v$ while the dot now follows the vector $v$. The result of the action of the operator  $u*v\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ on a state $w$ is then expressed as $u*v\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}w$. Here, $v\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}w$ represents the scalar product as defined above. Thus, the action of projection-type operators is self-explanatory. It follows from Eq. (5) that $(u*v\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}})^{\dagger}=v*u\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$.
Let us write down some simple relations employing the above expression for projection-type operators. If $\{/m/\}$ is an orthonormal basis, then the identity operator $I$ can be expressed as $I=\sum_{m}/m/*/m/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. (Of course, this expression is equivalent to the expansions of state vectors over the set /m//m/, such as $u=\sum_{m}/m/*c_{m}$, where $c_{m}=/m/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}u$.) Similarly, if $\nu/$ is, e.g., a complete set of continuous observables then $I$ can be written as $\int d\nu/\nu/*/\nu/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$.
Inserting projection operators into scalar products can, of course, be helpful. For example, using the latter expression for $I$, one obtains
$q/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}\psi\equiv/q/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}I\psi=\int d\nu/q/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}/\nu/*/\nu/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}\psi.$
In particular, the $q$-representation wave function is expressed in terms of the $\nu$-representation wave function in this way.
Let $\{/m/\}$ and $\{/n/\}$ be orthonormal bases. One may formally represent any operator $O$ in terms of the projection-type operators $m/*/n/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$:
$O=\sum_{m,n}\,/m/O/n/*/m/*/n/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}.$
In particular, one may write the identity operator $I$ as $\sum_{m,n}\,/m/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}/n/*/m/*/n/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$, which reduces to the expression for $I$ written above if the bases $\{/m/\}$ and $\{/n/\}$ coincide.
If $O_{pt}$ is a projection-type operator then the product $O_{1}*O_{pt}*O_{2}$ of three operators can be formally represented as a single projection-type operator. Denote this latter operator as $\mu/*/\nu/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ and $O_{pt}$ as $m/*/n/\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. Then we have $\mu/=O_{1}/m/$ and $\nu/=O_{2}^{\dagger}/n/$.
One may write the product of projection-type operators as such an operator times the product of scalar products, $\prod_{i=1}^{n}u_{i}*v_{i}\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}=\left(\prod_{i=1}^{n-1}v_{i}\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}u_{i+1}\right)u_{1}*v_{n}\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. All this shows the utility of the explicit expression for projection-type operators.
Up to this point, we formulated the present scheme as a one-space scheme. Let us incorporate the dual-space point of view in the present scheme. To this aim, we will reinterpret our notation. In the above scalar product expression $\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v, uu$ was considered a state vector, and the dot represented the scalar product sign. And now we will treat $u\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ as an indivisible symbol. Namely, we will consider $u\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ as the other designation for the linear functional $G_{u}$ in the relation $G_{u}(v)=u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$. That is, one will understand $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ as $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}(v)$ under this interpretation. Consequently, we will interpret $u\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ as a covector. This interpretation was proposed in the preliminary version of the present paper efr . [^7] In particular, it has been pointed out in efr (since the 2023 [v7] version) that an easy way to learn the bra-ket formalism is to proceed from this interpretation, as described below. A similar approach is also advocated in the recent paper lev . (The notation $Ou\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ has the same meaning as $(Ou)\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. In contrast, if $c$ is a constant then $cu\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ is interpreted as the product of the constant $c$ and the covector $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. If $c_{i}$ are constants and $u_{i}$ are state vectors, then we have $(\sum_{i}c_{i}u_{i})\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}=\sum_{i}c_{i}^{*}u_{i}\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. If, for example, $c_{i}$ and $c_{i}^{*}$ are scalar products, then for clarity they should be replaced by $c_{i}*$ and $c_{i}^{*}*$.)
In this way, one shifts to the dual-space description discussed in the previous section regarding the bra-ket formalism. According to this description, one deals with the spaces of state vectors and state covectors. The scalar product $u\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}v$ becomes the ”inner product” of a state covector $u\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ and a state vector $v$. And in both one-space and dual-space versions of the present scheme, the projection-type operator $u*v\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ is the ”outer product” of a state vector $u$ and a state covector $v\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$.
Thus, in contrast to the bra-ket, dual-space, formalism, the present formulation allows both the one-space and dual-space interpretations.
As mentioned in the previous section, the bra-ket expressions for covectors, such as $\langle\psi|O$ or $\langle\psi|O^{\dagger}$, entail potential ambiguities regarding the domain of definition. This drawback is absent here. Indeed, for instance, the covector $\langle\psi|O^{\dagger}$ is expressed as $O\psi\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ in the present formulation. From this expression, it is apparent that $\psi$ should belong to the domain of definition of the operator $O$.
(One can express a mapping $\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}\rightarrow O\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$ of covectors in terms of an operator acting on the covectors $\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}$. Let us denote this operator as $\bar{O}$ and position it to the right of covectors. Then, we have $O\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}\equiv\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}{\bar{O}}$. Accordingly, scalar products of the form $O\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}\chi$ can be rewritten as $(\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}{\bar{O}})\chi$. If the operator $O$ is linear, then the corresponding operator $\bar{O}$ is also linear. In terms of $\bar{O}$, the relation defining conjugate operators takes the form $(\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}{\bar{O}})\chi=\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}O^{\dagger}\chi$. If this relation holds for any states $\psi$ and $\chi$, one may view the operator $\bar{O}$ as the operator $O^{\dagger}$ acting to the left, i.e., on covectors. Accordingly, in this case, one can denote $\bar{O}$ as $O^{\dagger}$, assuming that in the expression $\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}O^{\dagger}\chi$ the operator $O^{\dagger}$ may act both to the right and to the left. This applies, in particular, in the finite-dimensional case.)
According to the above, the correspondence between the bra-ket notation and the notation of the dual-space version of the present scheme is as follows,
$|\psi\rangle\Leftrightarrow\psi,\qquad\langle\psi|\Leftrightarrow\psi\,\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}},\qquad\langle\psi|O\Leftrightarrow O^{\dagger}\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}.$
Further, when transitioning from the bra-ket formulation to the present formulation, the matrix element expression $\langle\psi|O|\chi\rangle$ should be replaced with either of the expressions $\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}O\chi$ and $O^{\dagger}\psi\mathbin{\stackrel{{\scriptstyle\centerdot}}{{}}}\chi$, provided that each of the latter expressions is well-defined. And if one of these expressions is not well-defined, then the bra-ket expression $\langle\psi|O|\chi\rangle$ has no sense, as discussed in Sec. 2 C. An easy opportunity to master the widely used bra-ket formalism is to start with the one-space version of the present scheme and then move on to its dual-space version, utilizing the listed correspondence.
IV Conclusion
To carry out representation-free considerations in quantum theory, the bra-ket formalism is generally employed. While it offers helpful tools for this purpose, it also has inherent drawbacks. Some of them arise from the fact that the bra-ket scheme is not well-suited for handling the domains of definitions of operators involved. One more drawback is that the bra-ket formalism, as a dual-space formalism, excludes the natural one-space point of view. Sec. 2 provides an analysis of the bra-ket formalism and addresses its drawbacks.
The present work gives a formulation of the representation-free approach which is entirely free of the drawbacks of the bra-ket formalism. In particular, unlike the bra-ket formalism, the present scheme allows both the needed one-space and dual-space interpretations. With all that, the present formulation provides efficient means for conducting representation-free considerations, akin to all such means offered by the bra-ket formalism. Herewith, the present scheme is less involved than the bra-ket one. The constructed ”universal” scheme is very well-suited for performing representation-free practical calculations.
Moving on from the bra-ket formalism to the present formulation is easy. And an easy opportunity to master the widely used bra-ket formalism would be to proceed from the present formulation.
V Acknowledgements
The author is grateful to A. Aiello and A. L. Barabanov, whose comments helped to improve the presentation. The author is indebted to F. Gieres for his helpful feedback regarding Ref. gie .